### v2

# Assessment of Spatio-Temporal Predictors in the Presence of Missing and Heterogeneous Data

Abstract
Deep learning approaches achieve outstanding predictive performance in modeling modern data, despite the increasing complexity and scale. However, evaluating the quality of predictive models becomes more challenging, as traditional statistical assumptions often no longer hold. In particular, spatio-temporal data exhibit dependencies across both time and space, often involving nonlinear dynamics, non-stationarities, and missing observations. As a result, advanced predictors such as spatio-temporal graph neural networks require novel evaluation methodologies. This paper introduces a residual correlation analysis framework designed to assess the optimality of spatio-temporal predictive neural models, particularly in scenarios with incomplete and heterogeneous data. By leveraging the principle that residual correlation indicates information not captured by the model, this framework serves as a powerful tool to identify and localize regions in space and time where model performance can be improved. A key advantage of the proposed approach is its ability to operate under minimal assumptions, enabling robust evaluation of deep learning models applied to multivariate time series, even in the presence of missing and heterogeneous data. The methodology employs tailored spatio-temporal graphs to encode sparse spatial and temporal dependencies within the data and utilizes asymptotically distribution-free summary statistics to pinpoint time intervals and spatial regions where the model underperforms. The effectiveness of the proposed residual analysis is demonstrated through validation on both synthetic and real-world scenarios involving state-of-the-art predictive models.
1 Introduction
Spatio-temporal predictive neural models fit data by leveraging the inductive biases inherent in relational information derived from the temporal and spatial dimensions of observations (Seo et al., 2018; Bai et al., 2020; Yu et al., 2018; Wu et al., 2019; Pal et al., 2021; Ruiz et al., 2020; Jin et al., 2023). The spatial domain is often represented as a graph, corresponding to structures such as a pixel grid, a 3D mesh, a road map, or a brain network (Shuman et al., 2013; Li et al., 2018; Stankovic et al., 2020). However, the term spatial should be interpreted more broadly, encompassing any functional dependencies among sensors, extending beyond correlations tied to their physical positions (Cini et al., 2023a). Modern spatio-temporal data, such as multivariate time series generated by sensor networks, present significant challenges, including irregular sampling, substantial missing observations, and the dynamic nature of heterogeneous sensors that can be added or removed over time (Alippi, 2014; Montero-Manso and Hyndman, 2021). In such contexts, predictive deep neural models may exhibit varying performance across different regions of the spatio-temporal domain, complicating the analysis of model quality and the detection of unexpected behaviors. Assessing the optimality of these predictors, particularly when leveraging advanced processing neural architectures, remains an intricate task for which no robust and effective methods currently exist.
The quality of deep predictive models is typically assessed using task-specific accuracy metrics, with the squared error being a common choice evaluating the 2-norm of the prediction residuals – the difference between the observed values and the model’s predictions; alternative metrics include absolute (MAE) and relative errors (MAPE) (Gneiting, 2011). These approaches are practical and straightforward, which likely explains their widespread adoption. However, they are inherently comparative, selecting the best model based on statistically superior performance relative to others. Consequently, they provide no direct insight into a model’s optimality or specific guidance on areas that require improvement.
An alternative approach to assessing model quality focuses on analyzing the correlations among prediction residuals rather than their magnitude. The rationale is that correlated residuals indicate structural information not captured by the model (Li, 2003), hence suggesting room for improvement. Over the years, various hypothesis tests have been developed to detect residual dependencies (Durbin and Watson, 1950; Ljung and Box, 1978; Box et al., 2015). Commonly known as randomness or whiteness tests, these tests assess whether residuals exhibit white noise behavior, meaning they lack correlations. However, they rely on strict assumptions, including fully available multivariate time series data, synchronous sampling, and identically distributed data. In practice, though, incomplete and heterogeneous data are the norm, posing significant challenges for existing methods and highlighting the need for more robust tests in such conditions (Zambon and Alippi, 2022).
1.1 Contributions
This paper presents a novel residual analysis framework for assessing the quality of models designed for spatio-temporal prediction tasks, particularly in scenarios with missing and heterogeneous data, and dynamic (spatial) relational structures. The framework not only detects residual correlations but also pinpoints specific regions where models fail to capture the underlying data-generating process, providing deeper insights beyond traditional evaluation methods. In particular, the paper addresses three key questions about model optimality, intended here as the absence of autocorrelations and cross-dependencies in prediction residuals.
  1. Q1
    Is the non-linear neural model optimal in terms of the absence of autocorrelations and cross-dependencies among prediction residuals?
  2. Q2
    Are there specific spatial regions (e.g., groups of time series) where predictions can be improved?
  3. Q3
    Are there specific time intervals where the model fails?
These questions are addressed by partitioning spatial and temporal regions of residuals into tailored subgraphs, followed by the computation of summary statistics, or scores, to quantify potential correlations. These scores measure spatio-temporal correlations by leveraging the AZ-whiteness test (Zambon and Alippi, 2022), which requires neither prior knowledge of the data distribution nor identical distributions across time series.
The key contributions of this paper are as follows:
  • A correlation scoring function that improves over the one introduced in (Zambon and Alippi, 2022), yielding scores that facilitate comparative analysis and interpretation.
  • A method to identify heterogeneous time series exposing missing data whose associated prediction residuals exhibit significant evidence of correlation.
  • A method to pinpoint time intervals where residuals display correlation.
  • A method to identify spatio-temporal regions where residual correlations are particularly prominent.
The AZ-analysis is validated using synthetic data, and its practical relevance is demonstrated in the real-world scenarios of traffic flow and energy production neural forecasting.
1.2 Significance and impact
Operating under minimal requirements, the proposed approach is versatile and broadly applicable to real-world scenarios involving deep and graph-based predictors. Notably, it does not impose assumptions on the distribution of residuals or require data to be identically distributed. The only prerequisite is that residuals are centered around zero – an assumption that is typically satisfied in most practical settings.
A key strength of the proposed analysis lies in its graph-based processing, which focuses on residual pairs that are more likely to exhibit correlations. This targeted approach enhances the statistical power of the method, making it more effective in detecting model shortcomings.
While the proposed residual analysis does not directly quantify the extent of potential model improvement, it complements traditional comparative evaluations based on prediction accuracy by providing an independent, metric-agnostic assessment of model quality. Experimental results reveal valuable insights from residual correlation analysis that are not captured through standard prediction error evaluations.
The paper is structured as follows. Section 2 reviews related work. Section 3 introduces the edge scores, the foundation of the proposed residual analysis, and shows how to design the spatio-temporal graph – a multiplex graph that links residual observations based on spatial and temporal relationships. Section 4 provides an overview of the AZ-whiteness test. The paper’s novel contributions are detailed in Section 5 and validated using synthetic data in Section 6. Sections 7 and 8 showcase two real-world applications. Finally, conclusions are presented in Section 9.
2 Related work
Statistical tests designed to discover correlations in temporal data date back to Durbin and Watson’s investigations (Durbin and Watson, 1950; Durbin, 1970) with variants inspecting longer-range correlations (Box and Pierce, 1970; Ljung and Box, 1978) and multivariate time series (Hosking, 1981; Li et al., 2019).
More general statistical tests leverage graph-based structures to connect data points. One of the earliest examples is the two-sample Friedman-Rafsky test (Friedman and Rafsky, 1979), which assesses the equality of distributions. Building on this foundation, subsequent research has extended these methods to change-point detection tests and statistical tests applicable to high-dimensional and non-Euclidean data (Chen and Zhang, 2013; Chu and Chen, 2019; Chen and Chu, 2023).
Following the approach in (Wald and Wolfowitz, 1940; Geary, 1970), which employs sign statistics to design non-parametric tests, Zambon and Alippi (2022) introduced a whiteness test to evaluate the overall presence of spatial and/or temporal correlation in prediction residuals.
To the best of our knowledge, this paper is the first one providing a comprehensive spatio-temporal analysis of correlation patterns at the node (time series), temporal, and local spatio-temporal levels.
Figure 1: Representation of spatio-temporal data $\boldsymbol{\mathbf{x}}$ as a set of time series with associated sequence of graphs $(\boldsymbol{\mathbf{g}}_{1},\boldsymbol{\mathbf{g}}_{2},\dots,\boldsymbol{\mathbf{g}}_{T})$ encoding functional relations. Observation $\boldsymbol{\mathbf{x}}_{\tau,v}$ at time step $\tau$ and node/sensor $v$ is multivariate. Nodes need not be available at all times (light gray boxes) and the graph topology can vary. $w_{t,e}$ denotes the weight of edge $e$ at time step $t$.
3 Preliminaries
3.1 Spatio-temporal predictions
Consider a multivariate time series
$\boldsymbol{\mathbf{x}}\doteq\{\boldsymbol{\mathbf{x}}_{t,v}:\ v\in V_{t},\;t=1,2,\dots\}$ (1)
where $\boldsymbol{\mathbf{x}}_{t,v}\in\mathbb{R}^{d_{x}}$ is a stochastic vector associated with observation of the $v$-th time series at time step $t$. We may refer to $v\in V_{t}$ as the sensor associated with the $v$-th time series from the sensor set $V_{t}$ available at time step $t$; denote $\boldsymbol{\mathbf{x}}_{t}\doteq\{\boldsymbol{\mathbf{x}}_{t,v}:v\in V_{t}\}$ and $V\doteq\bigcup_{t}V_{t}$. Consider a family of predictive models $f_{\theta}$ with parameter vector $\theta$ trained to predict at every time step $t$ target $\boldsymbol{\mathbf{y}}_{t,v}\in\mathbb{R}^{d_{y}}$, for all $v$, starting from a set of inputs in $\boldsymbol{\mathbf{x}}$; exogenous variables, such as those containing information about different types of sensors, are assumed encoded in $\boldsymbol{\mathbf{x}}$. A common example of a predictive task is time series forecasting (Cini et al., 2023a). In this context, $\boldsymbol{\mathbf{y}}_{t}$ corresponds to the window of subsequent observations $\boldsymbol{\mathbf{x}}_{t:t+H}$, therefore $d_{y}=d_{x}\times H$; notation $\boldsymbol{\mathbf{x}}_{a:b}$ indicates $\{\boldsymbol{\mathbf{x}}_{t}:a\leq t<b\}$. The presented formulation naturally handles missing data in $\boldsymbol{\mathbf{x}}$ using the sets $V_{t}$, which may not necessarily be equal to $V$. Moreover, no specific homogeneity assumption is made on the data, and the capability to support heterogeneous sensors will become evident from the discussion in Section 4.
In this paper, we consider the existence of functional dependencies among observations either available or extracted from the data itself (Cini et al., 2023c). We embed dependencies at time $t$ in graph
$\boldsymbol{\mathbf{g}}_{t}\doteq(V_{t},E_{t},\boldsymbol{\mathbf{w}}_{t},\boldsymbol{\mathbf{x}}_{t})$ (2)
whose edges $e\in E_{t}\subseteq V_{t}\times V_{t}$ represent (directed or undirected) binary relations between nodes in $V_{t}$ at time $t$. Each node $v\in V_{t}$ corresponds to time series observation $\boldsymbol{\mathbf{x}}_{t,v}\in\boldsymbol{\mathbf{x}}_{t}$. Nonnegative scalar weights $w_{t,e}\in\boldsymbol{\mathbf{w}}_{t}$ might be associated with each edge $e\in E_{t}$ to encode the strength of the relation, e.g., the physical distance of the two sensors or the capacity of a link in a transportation network. The result is a graph sequence
$(\boldsymbol{\mathbf{g}}_{1},\boldsymbol{\mathbf{g}}_{2},\dots,\boldsymbol{\mathbf{g}}_{t},\dots),$ (3)
where node set $V_{t}$, edge topology $E_{t}$ and edge weights $\boldsymbol{\mathbf{w}}_{t}$ can change over time. A visual representation of the set of time series $\boldsymbol{\mathbf{x}}$ and their relations in $\{\boldsymbol{\mathbf{g}}_{t}\}$ is provided in Figure LABEL:fig:st-data.
Model predictions $\hat{\boldsymbol{\mathbf{y}}}_{t}=f_{\hat{\theta}}(\boldsymbol{\mathbf{x}}_{t-W:t})$ are typically obtained from a sliding window of input data, defined as $\boldsymbol{\mathbf{x}}_{t-W:t}$, encompassing the past $W$ time steps. The parameter vector $\hat{\theta}$ is learned by minimizing a prediction loss function, such as mean squared error (MSE) or mean absolute error (MAE), between the predicted values $\hat{\boldsymbol{\mathbf{y}}}_{t,v}$ and the ground truth targets $\boldsymbol{\mathbf{y}}_{t,v}$. More broadly, predictions at time step $t$ can account for topological information: $\hat{\boldsymbol{\mathbf{y}}}_{t}=f_{\hat{\theta}}(\boldsymbol{\mathbf{g}}_{t-W:t})$, as done in graph-based models like spatio-temporal graph neural networks (STGNNs) (Cini et al., 2023a), as well as past predictions $\hat{\boldsymbol{\mathbf{y}}}_{t-W:t}$ and target values $\boldsymbol{\mathbf{y}}_{t-W:t}$. Define prediction residuals as
$\boldsymbol{\mathbf{r}}_{t,v}\doteq\boldsymbol{\mathbf{y}}_{t,v}-\hat{\boldsymbol{\mathbf{y}}}_{t,v}\ \in\mathbb{R}^{d_{y}}$ (4)
for all $v$ and $t$; the $i$-th component $\boldsymbol{\mathbf{r}}_{t,v}(i)$ of vector $\boldsymbol{\mathbf{r}}_{t,v}$ is the residual associated with the $i$-th target variable $\boldsymbol{\mathbf{y}}_{t,v}(i)$.
3.2 Spatio-temporal graphs
A finite sequence $(\boldsymbol{\mathbf{g}}_{1},\boldsymbol{\mathbf{g}}_{2},\dots,\boldsymbol{\mathbf{g}}_{T})$ of length $T$ and the associated set of residuals
$\boldsymbol{\mathbf{r}}\doteq\{\boldsymbol{\mathbf{r}}_{t,v}:v\in V_{t},t=1,\dots,T\}$ (5)
can be represented as a multiplex graph
$\boldsymbol{\mathbf{g}}^{*}\doteq(V^{*},E^{*},\boldsymbol{\mathbf{w}},\boldsymbol{\mathbf{r}})$ (6)
obtained by stacking all graphs $\boldsymbol{\mathbf{g}}_{t}$, for all $t=1,2,\dots,T$, and joining each node at time $t$ with itself at time $t-1$ as shown in Figure 2; in the figure, solid gray lines denote edges in $\boldsymbol{\mathbf{g}}_{t}$, while dashed gray lines indicate temporal edges across consecutive graphs. Moreover, $\boldsymbol{\mathbf{g}}^{*}$ is augmented with the residuals $\boldsymbol{\mathbf{r}}$ as node signals and a set $\boldsymbol{\mathbf{w}}$ of edge weights. As we will see in the sequel, the spatio-temporal graph $\boldsymbol{\mathbf{g}}^{*}$ provides a convenient static representation of the observed system dynamics.
Formally, node set $V^{*}$ is defined as
$V^{*}\doteq\left\{v_{t}\doteq(t,v):v\in V_{t},\ t=1,2,\dots,T\right\},$ (7)
where each node $v_{t}$ is associated with the residual vector $\boldsymbol{\mathbf{r}}_{t,v}\in\boldsymbol{\mathbf{r}}$. Edge set $E^{*}$ is the union of sets
$E^{*}\doteq E_{\rm sp}\cup E_{\rm tm}$ (8)
where
$E_{\rm sp}$$\doteq\left\{\{u_{t},v_{t}\}:(u,v)\in E_{t},v\neq u,1\leq t\leq T\right\},$ (9)
$E_{\rm tm}$$\doteq\left\{\{v_{t},v_{t+1}\}:v\in V_{t}\cap V_{t+1},1\leq t<T\right\}.$ (10)
$E_{\rm sp}$ collects all (spatial) edges in sets $E_{t}$, for all $t$, regardless of their orientation whenever it applies, while $E_{\rm tm}$ is the set of edges connecting corresponding nodes along the temporal dimension. Note that edge directions and self-loops are excluded in (9). However, this have no impact on the following analysis, as the statistics derived from equation (16) disregard directional information. Secondly, our focus is on the correlations between different observations in both time and space. Finally, if $\boldsymbol{\mathbf{g}}_{t}$ is undirected, weight $w_{u_{t},v_{t}}$ of spatial edge $\{u_{t},v_{t}\}\in E_{\rm sp}$ in $\boldsymbol{\mathbf{g}}^{*}$ is set equal to weight $w_{t,\{u,v\}}$ of $\boldsymbol{\mathbf{g}}_{t}$, i.e., the weight corresponding to edge $\{u,v\}\in E_{t}$. Conversely, for a directed graph $\boldsymbol{\mathbf{g}}_{t}$, $w_{u_{t},v_{t}}$ equals either $w_{t,(u,v)}$ or $w_{t,(v,u)}$, depending on which directional edge is in $E_{t}$, and $w_{u_{t},v_{t}}=w_{t,(u,v)}+w_{t,(v,u)}$ if both $(u,v)$ and $(v,u)$ are present. The weights for temporal edges can be defined arbitrarily or, e.g., to balance the overall impact of the spatial and temporal edges, as discussed in Section 4.
4 AZ-whiteness test
The analysis presented in this paper leverages the AZ-whiteness test (Zambon and Alippi, 2022), a statistical test designed to detect the presence of both autocorrelation and cross-correlation. The test is defined over hypotheses
$\begin{cases}H_{0}:&\text{All pairs }(\boldsymbol{\mathbf{r}}_{t,v},\boldsymbol{\mathbf{r}}_{\tau,u})\text{ are uncorrelated,}\\ H_{1}:&\text{At least a pair }(\boldsymbol{\mathbf{r}}_{t,v},\boldsymbol{\mathbf{r}}_{\tau,u})\text{ is correlated,}\end{cases}$ (11)
for $(t,v)\neq(\tau,u)$, and rejects the null hypothesis $H_{0}$ whenever the absolute value of statistic $C_{\lambda}(\boldsymbol{\mathbf{g}}^{*})$ exceeds a threshold $\gamma$ predetermined according to a user-defined significance level $\alpha$
$\mathbb{P}\left(\,|C_{\lambda}(\boldsymbol{\mathbf{g}}^{*})|\geq\gamma\mid H_{0}\,\right)=\alpha.$ (12)
The AZ-whiteness test statistic $C_{\lambda}(\boldsymbol{\mathbf{g}}^{*})\in\mathbb{R}$ is computed on the prediction residuals $\{\boldsymbol{\mathbf{r}}_{t,v}:\forall\ t,v\}$ with reference to the spatio-temporal multiplex graph $\boldsymbol{\mathbf{g}}^{*}$. More in detail, the test statistic is defined as
$C_{\lambda}(\boldsymbol{\mathbf{g}}^{*})\doteq\frac{\lambda\widetilde{C}_{\rm sp}+(1-\lambda)\widetilde{C}_{\rm tm}}{\left(\lambda^{2}W_{{\rm sp}}+(1-\lambda)^{2}W_{{\rm tm}}\right)^{1/2}},$ (13)
where parameter $\lambda\in[0,1]$ trades off spatial and temporal contributions
$\widetilde{C}_{\rm sp}$$\doteq\sum_{\{u_{t},v_{t}\}\in E_{\rm sp}}w_{\{u_{t},v_{t}\}}\;{\rm sgn}\left(\boldsymbol{\mathbf{r}}_{t,u}^{\top}\boldsymbol{\mathbf{r}}_{t,v}\right),$ (14)
$\widetilde{C}_{\rm tm}$$\doteq\sum_{\{v_{t},v_{t+1}\}\in E_{\rm tm}}w_{\rm tm}\;{\rm sgn}\left(\boldsymbol{\mathbf{r}}_{t,v}^{\top}\boldsymbol{\mathbf{r}}_{t+1,v}\right),$ (15)
respectively. Scalar quantities $\widetilde{C}_{\rm sp}$ and $\widetilde{C}_{\rm tm}$ are weighted sums assessing the spatial and temporal correlation, respectively, by accounting for positive and negative signs
${\rm sgn}\left(\boldsymbol{\mathbf{r}}_{t,v}^{\top}\boldsymbol{\mathbf{r}}_{\tau,u}\right)\ \in\{-1,0,1\}$ (16)
of the scalar product
$\boldsymbol{\mathbf{r}}_{t,v}^{\top}\boldsymbol{\mathbf{r}}_{\tau,u}=\sum_{i=1}^{d_{y}}r_{t,v}(i)\;r_{\tau,u}(i)$ (17)
between residual vectors $\boldsymbol{\mathbf{r}}_{t,v}$ and $\boldsymbol{\mathbf{r}}_{\tau,u}$; sign function ${\rm sgn}(a)$ equals $-1,0$ or $1$ depending on whether $a$ is negative, null, or positive. Scalar $w_{\rm tm}$ is a single weight assigned to all temporal edges and set such that the normalization term $W_{\rm tm}\doteq|E_{\rm tm}|\cdot w_{\rm tm}^{2}$ equals $W_{\rm sp}$
$W_{{\rm sp}}\doteq\sum_{\{u_{t},v_{t}\}\in E_{\rm sp}}w_{\{u_{t},v_{t}\}}^{2}.$ (18)
Figure 2: A section view of the spatio-temporal graph $\boldsymbol{\mathbf{g}}^{*}$ from Sec. 3.2. Each node is associated with a residual vector $\boldsymbol{\mathbf{r}}_{t,v}$ and each edge – either spatial (red) or temporal (blue) – is associated with a sign (16).
The distinction between spatial and temporal contributions is also highlighted in Figure 2. The intuition behind the test statistic (13) is that large positive values of $C_{\lambda}(\boldsymbol{\mathbf{g}}^{*})$ indicate the presence of adjacent residuals similarly oriented, thus suggesting correlation among variables; see also Figure 3 discussed below. Similarly, $C_{\lambda}(\boldsymbol{\mathbf{g}}^{*})\ll 0$ provides evidence of negative correlation.
Identifying a value of $\gamma$ for any choice of $\alpha$ is possible thanks to Theorem 1 showing that, asymptotically with the number of edges, $C_{\lambda}(\boldsymbol{\mathbf{g}}^{*})$ follows a standard Gaussian distribution irrespective of the distribution of the residuals and without requesting them to be identically distributed.
Theorem 1 (Zambon and Alippi (2022)).
Consider a spatio-temporal graph $\boldsymbol{\mathbf{g}}^{*}$ with associated stochastic residuals $\boldsymbol{\mathbf{r}}$ and the hyperparameter $\lambda\in[0,1]$. Assume
  1. A1
    All residual vectors $\boldsymbol{\mathbf{r}}_{t,v}$ in $\boldsymbol{\mathbf{r}}$ to be mutually independent and almost surely $\neq\boldsymbol{\mathbf{0}}$,
  2. A2
    $\mathbb{E}_{\boldsymbol{\mathbf{r}}_{t,v}}\left[{\rm sgn}\left(\bar{\boldsymbol{\mathbf{r}}}^{\top}\boldsymbol{\mathbf{r}}_{t,v}\right)\right]=\boldsymbol{\mathbf{0}}$ for all $\bar{\boldsymbol{\mathbf{r}}}\in\mathbb{R}^{d_{y}}\setminus\{\boldsymbol{\mathbf{0}}\}$, and $v_{t}\in V^{*}$,
  3. A3
    $w_{\rm tm},w_{t,\{u,v\}}\in[w_{-},w_{+}]$ for all $\{u_{t},v_{t}\}\in E_{t}$ and $t$, with $0<w_{-}<w_{+}$,
then, the distribution of $C_{\lambda}(\boldsymbol{\mathbf{g}}^{*})$ in (13) converges weakly to a standard Gaussian distribution $\mathcal{N}(0,1)$ as the number $|E^{*}|$ of edges goes to infinity.
The proof exploits the central limit theorem under the Lindeberg condition and is detailed in the original paper (Zambon and Alippi, 2022).
The assumptions of Theorem 1 are very mild, thanks also to the choice of assessing correlation by means of signs (16). More in detail, A1 is requested to meet the null hypothesis, while A2 and A3 are straightforwardly resolved in most scenarios, if not already fulfilled (Zambon and Alippi, 2022). In particular, for scalar residuals, A2 reduces to requesting that the residual median is null. Lastly, A3 requests positive bounded weights, as they are assumed here to encode the strength of the dependency between the related nodes. Whenever no weights are given, or they do not fulfill A3, we recommend setting all weights to $1$.
5 AZ-analysis of residuals
The AZ-analysis introduced in this paper addresses questions Q1Q3 by inspecting families $\{\boldsymbol{\mathbf{s}}_{i}\}$ of subsets $\boldsymbol{\mathbf{s}}_{i}$ of all possible pairs of residuals in $\boldsymbol{\mathbf{r}}$ and identifying those that display more prominent evidence of correlation according to scoring function $c_{\lambda}(\cdot)$, related to $C_{\lambda}(\cdot)$ as explained in next subsection. It is convenient to refer subsets $\boldsymbol{\mathbf{s}}_{i}$ as subgraphs of the fully connected graph connecting all pairs of observations in $\boldsymbol{\mathbf{r}}$. Given suitably designed subgraphs, we compute and compare correlation scores $c_{\lambda}(\boldsymbol{\mathbf{s}}_{i})$ for different subgraphs $\boldsymbol{\mathbf{s}}_{i}$, having in mind that $|c_{\lambda}(\boldsymbol{\mathbf{s}}_{i})|>|c_{\lambda}(\boldsymbol{\mathbf{s}}_{j})|$ implies the existence of a larger residual correlation in $\boldsymbol{\mathbf{s}}_{i}$ than in $\boldsymbol{\mathbf{s}}_{j}$. Although subgraphs $\{\boldsymbol{\mathbf{s}}_{i}\}$ may differ in size, we show that the associated scores $\{c_{\lambda}(\boldsymbol{\mathbf{s}}_{i})\}$ can be compared and ranked.
Figure 3: The figure compares the value of of statistics $C_{\lambda}(\cdot)$ defined in (13) (left) with that of the correlation scores $c_{\lambda}(\cdot)$ of (20) (right). Different levels of residual correlation and number $|E|$ of graph edges are considered. Each color corresponds to a different number of edges. Solid lines represent the expected value of the score estimated over 100 repeated experiments, dashed lines the standard deviation, shaded area the interquartile interval.
5.1 Correlation scores
Under the null hypothesis, test statistic $C_{\lambda}(\boldsymbol{\mathbf{s}})$ asymptotically follows a standard Gaussian distribution, regardless of the size of the graph $\boldsymbol{\mathbf{s}}$; see Theorem 1. This implies that all sample means should be centered around zero, except for those indicating evidence of correlation, while maintaining the same standard deviation, thereby ensuring a certain degree of comparability.
The left panel of Figure 3 illustrates values of $C_{\lambda}(\boldsymbol{\mathbf{s}})$ as function of the number of edges $|E_{\boldsymbol{\mathbf{s}}}|$ in $\boldsymbol{\mathbf{s}}$ and the Pearson correlation between residuals; detailed description of the experiment is provided in Appendix A. The figure illustrates that when no correlation exists among residuals, all scores are centered around zero with a standard deviation of 1, regardless of the number of edges considered. As correlation increases, each curve – corresponding to a different number of edges – edges—exhibits a monotonic increasing trend. However, we also observe that scores based on a larger number of edges are more effective in identifying correlated residuals. This suggests introducing a statistic that enables the comparison of score values independent of the number of edges used in their computation.
The maximum value of each curve is $\sqrt{|E_{\boldsymbol{\mathbf{s}}}|}$. In fact, when all signs in (13) equal 1, the test statistic [^1] For brevity, here $\boldsymbol{\mathbf{s}}$ is a generic graph with no explicit distinction between spatial and temporal edges. Accordingly, $\widetilde{C}$ is the weighted sum across all edges in set $E_{\boldsymbol{\mathbf{s}}}$ (regardless of whether they are spatial or temporal), $\lambda$ is incorporated in the weights $\boldsymbol{\mathbf{w}}$ and the denominator of (13) reduces to $\lVert\boldsymbol{\mathbf{w}}\rVert_{2}$. is
$C_{\lambda}(\boldsymbol{\mathbf{s}})=\frac{\widetilde{C}}{\sqrt{\lVert\boldsymbol{\mathbf{w}}\rVert_{2}^{2}}}=\frac{\lVert\boldsymbol{\mathbf{w}}\rVert_{1}}{\lVert\boldsymbol{\mathbf{w}}\rVert_{2}};$ (19)
note that in the experiment yielding to Figure 3 all edge weights are equal to 1, therefore $\lVert\boldsymbol{\mathbf{w}}\rVert_{1}=|E_{\boldsymbol{\mathbf{s}}}|$ and $\lVert\boldsymbol{\mathbf{w}}\rVert_{2}=\sqrt{|E_{\boldsymbol{\mathbf{s}}}|}$. Following this observation, we scale statistic $C_{\lambda}(\cdot)$ to yield values in the $[-1,1]$ interval as
$c_{\lambda}(\boldsymbol{\mathbf{s}})\doteq C_{\lambda}(\boldsymbol{\mathbf{s}})\frac{\lVert\boldsymbol{\mathbf{w}}\rVert_{2}}{\lVert\boldsymbol{\mathbf{w}}\rVert_{1}}=\frac{\widetilde{C}}{\lVert\boldsymbol{\mathbf{w}}\rVert_{1}},$ (20)
and name $c_{\lambda}(\cdot)$ correlation score.
The right panel of Figure 3 shows the behavior of scores $c_{\lambda}(\cdot)$. While they share the same advantages as $C_{\lambda}(\cdot)$, where magnitude increases as correlation increases, we also observe that their expected value remains the same regardless of the number of edges. Accordingly, we suggest to inspect those subgraphs $\boldsymbol{\mathbf{s}}_{i}$ associated with $|C_{\lambda}(\boldsymbol{\mathbf{s}}_{i})|\gg 0$ (evidencing correlation), starting from those with large $|c_{\lambda}(\boldsymbol{\mathbf{s}}_{i})|$ in decreasing order.
5.2 Subgraphs as sets of edges
Note that the number of correlated pairs of residuals is, in principle, quadratic in the number of residuals $\boldsymbol{\mathbf{r}}$ in (5), $N\,T$, where $T$ is the length of the time series and $N$ the average number of available sensors. More precisely, there are $N\,T$ observed residuals – corresponding to the nodes $v_{t}\in V^{*}$ of $\boldsymbol{\mathbf{g}}^{*}$ – which leads to $(N\,T)(N\,T-1)/2$ (possibly) correlated node pairs. Indeed, considering all possible subsets of node pairs is unfeasible, as their number is about $2^{{N^{2}\,T^{2}/2}}$. We deal with this complexity problem in two steps.
First, we confine the analysis to node pairs corresponding to edges in $\boldsymbol{\mathbf{g}}^{*}$, which amounts to considering all possible subgraphs of $\boldsymbol{\mathbf{g}}^{*}$. As the relational information encoded in multiplex $\boldsymbol{\mathbf{g}}^{*}$ is relevant to solving the downstream task, the node pairs are those that are more likely to exhibit correlation in the associated residuals. The inspection reduces to $2^{E_{*}}$ subgraphs, i.e., approximately $2^{N\,\delta\,T/2}$, with $\delta$ the average node degree in the graph sequence incremented by 2 to account for temporal edges. Second, as computing and inspecting scores for all $2^{E_{*}}$ subsets is still impractical, we propose principled families of subgraphs designed to address the three key questions Q1, Q2, and Q3 in Sections 5.3 to 5.6.
Figure 4: Top) Generic subgraphs $\boldsymbol{\mathbf{s}}_{v},\boldsymbol{\mathbf{s}}_{u}$ involved in the computation of node scores $c_{\lambda}(v)$ and $c_{\lambda}(u)$. Center) Generic subgraphs $\boldsymbol{\mathbf{s}}_{t},\boldsymbol{\mathbf{s}}_{\tau}$ involved in the computation of temporal scores $c_{\lambda}(t)$ and $c_{\lambda}(\tau)$. Bottom) Generic subgraphs $\boldsymbol{\mathbf{s}}_{t,v},\boldsymbol{\mathbf{s}}_{\tau,u}$ involved in the computation of temporal scores $c_{\lambda}(v,t)$ and $c_{\lambda}(\tau,u)$.
5.3 Determining the presence of correlation (Q1)
The AZ-whiteness test (13) with $\lambda=1/2$ applied to multiplex graph $\boldsymbol{\mathbf{g}}^{*}$ allows us for assessing the overall presence of correlation; here, large values of $|C_{\lambda=1/2}(\boldsymbol{\mathbf{g}}^{*})|$ denote a pronounced correlation in the residuals.
Instead, the values of statistics $C_{\lambda}(\boldsymbol{\mathbf{g}}^{*})$ with $\lambda=0$ and $\lambda=1$ provide first insights about the presence of temporal and spatial correlations, respectively. This follows from Theorem 1, as $C_{\lambda=0}(\boldsymbol{\mathbf{g}}^{*})$ and $C_{\lambda=1}(\boldsymbol{\mathbf{g}}^{*})$ share the same distribution under the assumption of independent residuals; therefore comparing their values indicates which – temporal or spatial – elements in the processing pipeline show margins for improvement.
For instance, $|C_{\lambda=0}(\boldsymbol{\mathbf{g}}^{*})|>|C_{\lambda=1}(\boldsymbol{\mathbf{g}}^{*})|$ with $|C_{\lambda=0}(\boldsymbol{\mathbf{g}}^{*})|\gg 0$ would indicate temporal (auto)correlation to be more problematic than spatial correlation. Should this be the case, the designer has to intervene on the temporal processing pipeline to improve the model, e.g., by changing the width of the input window or the neural prediction architecture; we refer to (Gao and Ribeiro, 2022; Cini et al., 2023a) for discussions about different ways of combining spatial and temporal processing in deep architectures.
5.4 Node-level analysis (Q2)
In order to study residuals associated with a subset $\phi\subset V$ of nodes/sensors, we propose node score
$c_{\lambda}(\phi)\doteq c_{\lambda}(\boldsymbol{\mathbf{s}}_{\phi})$ (21)
evaluated on subgraph $\boldsymbol{\mathbf{s}}_{\phi}$ extracted from multiplex $\boldsymbol{\mathbf{g}}^{*}=(V^{*},E^{*},\boldsymbol{\mathbf{w}},\boldsymbol{\mathbf{r}})$. Differently from Section 5.3, node score $c_{\lambda}(\phi)$ is a correlation score (20), not a statistic (13); this particular choice for the assessment function follows from the discussion on the score comparability of Section 5.1.
Subgraph $\boldsymbol{\mathbf{s}}_{\phi}$ is defined by all edges in $E^{*}$ with at least one ending node related to $\phi$, i.e., by all nodes from $V^{*}$ in set
$V_{\phi}^{*}\doteq\{v_{t}:v\in\phi\cap V_{t},\ t=1,\dots,T\}$ (22)
or the associated 1-hop neighborhood
$N(V_{\phi}^{*})\doteq\bigcup_{v_{t}\in V_{\phi}^{*}}N(v_{t});$ (23)
$N(v_{t})$ denotes the set of 1-hop neighbors of node $v_{t}$. Edges in $\boldsymbol{\mathbf{g}}^{*}$ linking nodes in $N(V_{\phi}^{*})\setminus V_{\phi}^{*}$ are excluded, as they do not directly impact any node of interest $v\in\phi$. Special cases analyzing single nodes $v$ derive from (21) by setting $\phi=\{v\}$, as shown at the top of Figure 4; with little abuse of notation, denote $c_{\lambda}(v)\doteq c_{\lambda}(\{v\})$.
Relevant considerations follow from inspecting $c_{\lambda}(v)$. For example, observing large values of $|c_{\lambda}(v)|$ when relying on global predictive models, like standard graph neural networks sharing parameters across nodes, can be a warning on the presence of local effects that would be better described by local, node-specific models (Montero-Manso and Hyndman, 2021; Cini et al., 2023b).
5.5 Temporal analysis (Q3)
Request Q3 can be addressed by considering subgraphs $\{\boldsymbol{\mathbf{s}}_{t}\}_{t}$ that slice multiplex graph $\boldsymbol{\mathbf{g}}^{*}$ along the temporal axis. For each time step $t$, define time score
$c_{\lambda}(t)\doteq c_{\lambda}(\boldsymbol{\mathbf{s}}_{t})$ (24)
determined by the subgraph $\boldsymbol{\mathbf{s}}_{t}$ corresponding to the nodes at time steps $t-1$, $t$, and $t+1$. More specifically, $\boldsymbol{\mathbf{s}}_{t}$ is the subgraph of $\boldsymbol{\mathbf{g}}^{*}$ defined by those edges with at least one ending node in
$V_{t}^{*}\doteq\{v_{t}:v\in V_{t}\},$ (25)
as depicted at the center of Figure 4.
Similarly to $c_{\lambda}(\phi)$ in (21), consider a time window $\omega=\{t_{1},t_{1}+1,\dots,t_{2}\}$, for some $t_{1}<t_{2}$, and construct subgraphs $\boldsymbol{\mathbf{s}}_{\omega}$ from
$V_{\omega}^{*}\doteq\{v_{t}:v\in V_{t},\ t\in\omega\},$ (26)
to define statistic
$c_{\lambda}(\omega)\doteq C_{\lambda}(\boldsymbol{\mathbf{s}}_{\omega}).$ (27)
Discovering time periods associated with large $|c_{\lambda}(\cdot)|$ might even indicate non-stationary behaviors that the given model is not able to properly capture. A strategy to deal with these time-variant processes would consider, e.g., online adaptive mechanisms updating the prediction model over time (Ditzler et al., 2015).
5.6 Local spatio-temporal analysis
As the number of nodes and time steps increase, scores (21) and (27) become less effective in discovering correlation existing in confined regions of the space-time; this is a consequence of $c_{\lambda}(\phi)$ and $c_{\lambda}(\omega)$ being averages across all nodes and all time steps, respectively. To address this limitation, we present a strategy to identify local correlation patterns across time and space altogether.
For each node $v_{t}\in V^{*}$ of $\boldsymbol{\mathbf{g}}^{*}$ associated with sensor $v$ at time step $t$ define local score
$c_{\lambda}(t,v)\doteq c_{\lambda}(\boldsymbol{\mathbf{s}}_{t,v}),$ (28)
as the correlation score evaluated on subgraph $\boldsymbol{\mathbf{s}}_{t,v}$ constructed from the $k$-hop neighbors of $v_{t}$, for some positive $k$; see the bottom part of Figure 4. Correlation score $c_{\lambda}(t,v)$ is local in the space-time and accounts for the correlation around $(t,v)$. In the following experiments, spatio-temporal scores $\{c_{\lambda}(t,v),v\in V,t=1,\dots,T\}$ are constructed by considering $k=4$ hops around each space-time location $(t,v)$.
More intricate correlation patterns can be highlighted by local scores $c_{\lambda}(t,v)$. An example is provided in Section 7 where data preprocessing issues have been evidenced.
AZ-test statistics: $\quad C_{\lambda=1}(\boldsymbol{\mathbf{g}}^{*})=16.3$, $\quad C_{\lambda=0}(\boldsymbol{\mathbf{g}}^{*})=18.9$, $\quad C_{\lambda=1/2}(\boldsymbol{\mathbf{g}}^{*})=24.3$. Refer to captionRefer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Scores involved in the analysis of residuals on synthetic data. Scores associated with $\lambda=0$, $1/2$, and $1$ are depicted in blue, green, and red colors, respectively. Top left) Temporal scores $c_{\lambda}(t)$; a moving average is applied to improve readability. Center) Node-level scores $c_{\lambda}(v)$ as both line plots and heatmaps on the graph. Bottom left) Local spatio-temporal score $c_{\lambda}(t,v)$ for $\lambda=1/2$; see Figure 6 for $\lambda=0$ and $1$. Red and blue boxes (sets $A$ and $B$) highlight regions with spatial and temporal correlation, respectively. Values of the AZ-test statistics are reported at the top left of the figure.
Figure 6: Synthetic residuals generated according to Section 6 (left). Scores $c_{\lambda=0}(t,v)$ (center) and $c_{\lambda=0}(t,v)$ (right).
AZ-test statistic: $\quad C_{\lambda=1}(\boldsymbol{\mathbf{g}}^{*})=14.6$, $\quad C_{\lambda=0}(\boldsymbol{\mathbf{g}}^{*})=13.9$, $\quad C_{\lambda=1/2}(\boldsymbol{\mathbf{g}}^{*})=20.2$. Refer to caption
Figure 7: Scores involved in the AZ-analysis of residuals on synthetic residuals with missing observations. White areas in the bottom-left figure represent regions of missing observations.
AZ-test statistic: $\quad C_{\lambda=1}(\boldsymbol{\mathbf{g}}^{*})=14.8$, $\quad C_{\lambda=0}(\boldsymbol{\mathbf{g}}^{*})=15.5$, $\quad C_{\lambda=1/2}(\boldsymbol{\mathbf{g}}^{*})=21.4$. Refer to caption
Figure 8: Scores involved in the AZ-analysis of residuals on synthetic residuals with heterogeneous. Residuals at different space-time locations are generated from different distributions.
6 Empirical validation of the scores behavior
The first experiment considers a synthetic dataset where we artificially induced temporal and spatial correlation to show that the proposed scores $c_{\lambda}(\cdot)$ are indeed able to detect correlation patterns.
6.1 Synthetic residuals
The residual signal $\boldsymbol{\mathbf{r}}$ is modeled as Gaussian noise exhibiting spatial and temporal correlations within specific regions defined by the intersection of sets $A$ and $B$, respectively. Outside $A\cup B$ residuals are i.i.d..
The data is generated for 60 nodes and 400 time steps. Set $A$ covers time steps in $[200,400)$ and nodes in $[15,45)$, whereas set $B$ time steps in $[100,300)$ and nodes in $[30,60)$. The considered graph topology and sets $A$ and $B$ are depicted in Figure 5. Starting from i.i.d. white noise $\varepsilon_{t,v}\sim\mathcal{N}(0,1)$, correlation is induced as follows. Within region $A\setminus B$ of only spatial correlation, residuals are obtained as
$r_{t,v}=\varepsilon_{t,v}+\frac{\sum_{u\in N(v)}\varepsilon_{t,u}}{|N(v)|},$ (29)
while in $B\setminus A$ only temporal correlation is present:
$r_{t,v}=\varepsilon_{t,v}+\frac{\varepsilon_{t-1,u}+\varepsilon_{t+1,u}}{2}.$ (30)
Both spatial and temporal correlations are introduced in $A\cap B$:
$r_{t,v}=\varepsilon_{t,v}+\frac{\sum_{u\in N(v)}\varepsilon_{t,u}}{|N(v)|}+\frac{\varepsilon_{t-1,u}+\varepsilon_{t+1,u}}{2}.$ (31)
Finally, for each point $(t,v)$ outside $A\cup B$, uncorrelated residuals $r_{t,v}$ equal $\varepsilon_{t,v}$. An example of correlated residuals generated from this process is depicted in the left part of Figure 6.
6.2 Analysis of residuals
Figure 5 provides a visual representation of the proposed analysis of residuals. We observe the following.
Overall presence of correlation
The values of the AZ-whiteness test statistics $C_{\lambda}(\boldsymbol{\mathbf{g}}^{*})$ are substantially larger than 00, correctly identifying the overall presence of both spatial and temporal correlation.
Spatial correlation
The magnitude of all node-level scores $c_{\lambda=1}(v)$ and temporal scores $c_{\lambda=1}(t)$ (red dot-dashed lines) are consistent with set $A$ containing spatial correlation.
Temporal correlation
Similarly, $c_{\lambda=0}(v)$ and $c_{\lambda=0}(t)$ (blue dashed lines) are consistent with $B$ associated with temporal correlation.
Spatio-temporal scores
$c_{\lambda=1/2}(v)$ and $c_{\lambda=1/2}(t)$ (green solid lines) are consistent with the union set $A\cup B$; local scores $c_{\lambda=1/2}(t,v)$ have similar behavior. Moreover, the scores are larger where both types of correlation are present ($A\cap B$), than where there is only either of the two ($A\setminus B$ and $B\setminus A$).
Also the values of the local spatio-temporal scores $c_{\lambda}(t,v)$ with $\lambda=1$ and $\lambda=0$ are consistent with both sets $A$ and $B$; we refer the reader to Figure 6.
6.3 Analysis with missing and heterogeneous data
Lastly, we report the scores of the residual analysis carried out on similar synthetic data where, however, about 20% of the observations were missing and where the data comes from heterogeneous sensors. For the scores of these experiments, we refer the reader to Figures 7 and 8.
Analysis with missing data
Despite the presence of data gaps, visually represented as white areas within the representation of $c_{\lambda=1/2}(t,v)$ in Figure 7, the reported correlation scores exhibit patterns aligned with those present in the data and discussed in previous Section 6.2.
To validate the effectiveness of the proposed analysis on data coming from heterogeneous sensors, the residuals $\boldsymbol{\mathbf{r}}$ are generated following the process described in Section 6.1, but starting from non-i.i.d. noise $\varepsilon$. In particular, each noise component $\varepsilon_{t,v}$ is sampled from either of three zero-median distributions: a uniform distribution, a Laplace distribution, a bimodal distribution generated from a mixture of 2 Gaussian distributions; the distribution assignment is uniform but random.
Analysis with heterogeneous data
The analysis reported in Figure 8 shows scores very similar to those of Figures 5 and 7. In particular, both regions $A$ and $B$ are clearly identifiable and analog conclusions can be drawn.
To conclude, the reported results validate the proposed AZ-analysis by showing that both spatial and temporal correlation can be identified. In particular, it can be identified at a global level (Q1), at the level of single nodes (Q2) and of time steps (Q3). Secondly, the results confirm the broad applicability of the presented residual analysis which can effectively operate under mild assumptions, and in presence of missing and heterogeneous data.
7 Use case in traffic forecasting
In the second experiment, we analyze the prediction residuals of spatio-temporal graph neural networks (STGNNs) trained for multi-step ahead forecasting on traffic flow data.
7.1 Data and models
The study utilizes the MetrLA traffic dataset (Li et al., 2018), which aggregates traffic information into 5-minute intervals of sensor readings from March to June 2012. The dataset comprises 207 univariate time series. A graph connecting the sensors is constructed from their pairwise distance. The provided data contain missing observations (approximately 8% of the data), which were imputed using preceding observations.
The selected set of predictive models includes generic baselines for time series forecasting and well-established methods from the literature: GWNet (Wu et al., 2019), DCRNN (Li et al., 2018), and AGCRN (Bai et al., 2020). The baseline models consist of a recurrent neural network (RNN) with gated recurrent units (GRUs) trained and operated on individual univariate time series (uvRNN), a second RNN trained and operated on the multivariate time series obtained by stacking all univariate time series (mvRNN), and a generic time-then-space STGNN (ttsRNN) processing the output of uvRNN with a graph neural network (Cini et al., 2023b). Additional experimental details are provided in Appendices B and C.
(a) Multi-step ahead forecasting
Score $c_{\lambda=1/2}(t)$ Refer to caption
MAE Refer to caption
(b) 1-step ahead forecasting
Score $c_{\lambda=1/2}(t)$ Refer to caption
Figure 9: Left) Temporal scores $c_{\lambda=1/2}(t)$ of the residuals (top) and predictive MAE (bottom) on MetrLA. Shaded areas highlight regions of imputed target data. Right) Temporal scores $c_{\lambda=1/2}(t)$ on 1-step ahead prediction on MetrLA. Shaded areas highlight regions of imputed target data.
Figure 10: Extended residual analysis performed on MetrLA for a single model. Regions of the imputed data are reported as dark regions in the plot on the right.
7.2 Collected insights and discussion
Figure 9(a) presents the time scores $c_{\lambda=1/2}(t)$ and prediction error (expressed here in terms of MAE) for all considered models. For visualization purposes, a window of 500 time steps and 40 nodes is depicted in the figures.
Identification of correlation patterns
From the top part of Figure 9(a), we observe that regardless of the predictive model, the time scores are all consistently high within two timeframes around $t=100$ and $t=450$.
As indicated by the shaded areas in the figure, these regions correspond to time steps with artificially imputed data. The observed high scores likely relate with the imputation method employed, which replicates the last observed value.
Supplementing performance-based analyses
The bottom part of Figure 9(a) displays the time-wise MAE of all predictors. Comparing the MAE with scores $c_{\lambda=1/2}(t)$, we observe that the higher correlation in the shaded areas is not accompanied by a significant increase in prediction error. This finding demonstrates that the proposed residual analysis can reveal valuable insights that are not apparent from analyses solely based on prediction error
Predictive horizon
Comparing Figure 9(a), which computes time scores from multi-step ahead predictions, with Figure 9(b) focusing on 1-step ahead predictions, reveals that the correlation patterns are more pronounced in the latter case. Furthermore, the observed correlation is diminished in regions where data is not imputed, suggesting that there is less room for improvement in short-term predictions compared to long-term predictions.
Extending the residual analysis for one of the models, GWNet, yields further relevant insights; please, refer to Figure 10.
Local correlation patterns
Scores $c_{\lambda}(t,v)$ confirm that the aforementioned timeframes are indeed problematic. Moreover, they highlight specific nodes, namely nodes 28 and 29, that warrant closer attention. Further examination of the mask of imputed data at the bottom of Figure 10 reveals that the time series for these nodes have also been imputed.
Figure 11: Analysis of the prediction residuals for energy production prediction. Top) Time scores $c_{\lambda}(t)$ and MAE of three days of energy production. Bottom) Relative prediction error (MAPE) and mean absolute value of the target $\boldsymbol{\mathbf{y}}_{t}$. Absent observations correspond to the null solar radiation during nighttime. Shaded area denote the interquartile range.
8 Use case in energy production
In the third experiment, we investigate the task of forecasting energy production from photovoltaic plants.
8.1 Data and models
We consider EngRAD dataset (Marisca et al., 2024), which encompasses 5 meteorological variables related to photovoltaic energy production. Hourly estimations of these variables are obtained from a grid of 487 locations across England. A graph representation is constructed based on the pairwise geographical distance between locations. The ttsRNN model was employed and trained to minimize the MAE on 3-hour ahead forecasts of horizontal irradiance, leveraging an input window of 24 preceding hours and related weather variables. The residual analysis was conducted on the test set, focusing on 1-hour ahead forecasts. Further experimental details are provided in Appendices B and C.
8.2 Insights from the residual analysis
Figure 11 reveals a daily trend evident in both the MAE and the time scores $c_{\lambda}(t)$, attributable to the natural solar irradiation cycle. Missing observations correspond to the negligible solar irradiation of the night.
Correlation at dawn and dusk
Higher correlations observed during dawn and dusk suggest room for model improvement, while the MAE remains relatively low throughout these transitional periods. While the MAE alone is not particularly concerning during dawn and dusk, further inspection reveals higher relative prediction error (mean absolute percentage error, MAPE), reinforcing the evidence provided by $c_{\lambda}(t)$ that the model’s predictions can improve in these hours of the day.
Furthermore, the residual analysis unveils another interesting insight: in certain instances where the model exhibits low MAE values, a high correlation is present, thus indicating that they can be further enhanced. Specifically, on June 23rd and 24th, 2020, the target variable exhibited a notably high magnitude with reduced variability, coupled with a consistently lower MAE. While this suggests that the target variable is easier to forecast during these specific periods, it is likely that the training primarily focused on regions characterized by higher prediction errors.
9 Conclusions
In this paper, we introduce a novel analysis of model prediction residuals, called AZ-analysis, for spatio-temporal data that allows practitioners to identify space-time regions, such as time intervals or sets of sensors, where the model performance can be improved. The proposed analysis complements that of the prediction error by evaluating the presence of temporal and spatial residual correlation in regions of the space-time and sheds light on aspects where the predictive model is not optimal. The AZ-analysis is well-suited for assessing deep neural models in real-world scenarios as (i) it can operate in presence of missing heterogeneous data, (ii) it scales to large and complex datasets by exploiting graph-side information, and (iii) only requests residuals centered around zero in a few setups; no other assumption about the distribution of residuals is requested, not even of identical distribution. The proposed framework is first validated on synthetic experiments and general comments on how to interpret the results are provided along with real-world use cases in the context of forecasting traffic flow and energy production to showcase the potential of the proposed AZ-analysis.
Appendix A Experiment on the comparability of scores
In the experiment of Figure 3, Section 5.1, the values of statistic $C_{\lambda}(\boldsymbol{\mathbf{s}})$ and score $c_{\lambda}(\boldsymbol{\mathbf{s}})$ are computed for graphs $\boldsymbol{\mathbf{s}}$ of different number of edges $|E_{\boldsymbol{\mathbf{s}}}|$ and Pearson correlation among residuals. Considered subgraph $\boldsymbol{\mathbf{s}}$ is given as a collection of non-incident edges. For each edge $(u,v)$, residual $\boldsymbol{\mathbf{r}}_{v}\in\mathbb{R}^{d}$ of node $v$ is generated as
$\boldsymbol{\mathbf{r}}_{v}=\rho\,\boldsymbol{\mathbf{r}}_{u}+\sqrt{1-\rho^{2}}\,\boldsymbol{\mathbf{r}}_{u}^{\prime}+\boldsymbol{\mathbf{m}},$ (32)
where $\boldsymbol{\mathbf{r}}_{u}$ is the residual at node $u$. The components of vectors $\boldsymbol{\mathbf{r}}_{u}$ and $\boldsymbol{\mathbf{r}}_{u}^{\prime}$ are independent samples from the standard Gaussian distribution, whereas $\boldsymbol{\mathbf{m}}$ centers the median of $\boldsymbol{\mathbf{r}}_{v}$ to zero. Note that scalar $\rho$ is selected in the $[0,1]$ interval and determines the Pearson correlation between $\boldsymbol{\mathbf{r}}_{v}$ and $\boldsymbol{\mathbf{r}}_{u}$. In fact, as $\mathbb{E}[\boldsymbol{\mathbf{r}}_{u}]=\mathbf{0}$, ${\rm Var}[\boldsymbol{\mathbf{r}}_{u}]=\mathbb{I}$ and $\mathbb{E}[\boldsymbol{\mathbf{r}}_{v}]=\boldsymbol{\mathbf{m}}$ by construction,
${\rm Var}[\boldsymbol{\mathbf{r}}_{v}]=\rho^{2}{\rm Var}[\boldsymbol{\mathbf{r}}_{u}]+(1-\rho^{2}){\rm Var}[\boldsymbol{\mathbf{r}}_{u}^{\prime}]=\mathbb{I}$ (33)
and
$\text{Cor}[\boldsymbol{\mathbf{r}}_{u},\boldsymbol{\mathbf{r}}_{v}]$ $=\mathbb{E}[(\boldsymbol{\mathbf{r}}_{v}-\boldsymbol{\mathbf{m}})\boldsymbol{\mathbf{r}}_{u}^{\top}]=$ (34)
$=\mathbb{E}[(\rho\,\boldsymbol{\mathbf{r}}_{u}+\sqrt{1-\rho^{2}}\,\boldsymbol{\mathbf{r}}_{u}^{\prime})\boldsymbol{\mathbf{r}}_{u}^{\top}]$ (35)
$=\rho\mathbb{E}[\boldsymbol{\mathbf{r}}_{u}\boldsymbol{\mathbf{r}}_{u}^{\top}]+\sqrt{1-\rho^{2}}\mathbb{E}[\boldsymbol{\mathbf{r}}_{u}^{\prime}\boldsymbol{\mathbf{r}}_{u}^{\top}]$ (36)
$=\rho\,\mathbb{I}+\sqrt{1-\rho^{2}}\mathbb{E}[\boldsymbol{\mathbf{r}}_{u}^{\prime}]^{\top}\mathbb{E}[\boldsymbol{\mathbf{r}}_{u}]=\rho\,\mathbb{I}.$ (37)
Moreover, above construction grants Assumption A2 to hold.
Dimension $d$ of residual vectors is set to 5. Mean, standard deviation, and 25% and 75% percentiles are reported in Figure 3 and are estimated over 100 repeated simulations.
Appendix B Additional details on models and training
For the traffic forecasting task, predictions are made from an input window of 12 time steps (1 hour), with the hour of the day and the day of the week added as covariates to all models. For the solar irradiance prediction, forecasts are issued for 6 hours ahead with an input window of 24 hours; here, the hour of the day and hour of the year are added as covariates. The residual analyses are conducted on the test set considering traffic forecasts at 5, 20, 40 and 60 minutes and solar irradiance predictions in the next hour for the energy production.
Baseline models uvRNN, mvRNN and ttsRNN are 1-layer RNN with gated recurrent units and the exponential linear unit as activation function. Each model has 32 hidden neurons. ttsRNN implements a simple message passing based on average aggregation on the uvRNN outputs. Models are trained to minimize the mean absolute error (MAE) on the first 70% of data, validated on the following 10%, and tested on the remaining 20%; missing and imputed data are appropriately masked out during training. Training is performed for 200 epochs using Adam optimizer with an initial learning rate of 0.003 decreased every 50 epochs by a factor of 4. Each epoch consists of 300 batches of size 64. Early stopping is triggered after 50 epochs of no improvement on the validation set.
Appendix C Hardware and software
Model training and inference are performed on a workstation equipped with Intel(R) Xeon(R) Silver 4116 CPU 2.10GHz processors and NVIDIA TITAN V GPUs. The analysis of residuals does not require GPU acceleration. Experiments are developed in Python 3.8 mainly relying on open-source libraries PyTorch (Paszke et al., 2019), PyTorch Geometric (Fey and Lenssen, 2019), Torch Spatiotemporal (Cini and Marisca, 2022) and NumPy (Harris et al., 2020). The source code to reproduce the results will be released upon publication.
Acknowledgements
This work was supported by the Swiss National Science Foundation project FNS 204061: High-Order Relations and Dynamics in Graph Neural Networks.